an equation of a line Containing the point (2,5)with the given Slope of m=5
Answer by lwsshak3(6522)

an equation of a line Containing the point (2,5)with the given Slope of m=5
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept.
y=5x+b
solve for b using given point (2,5)
5=5*2+b
b=5-10=-5
Equation:
y=5x-5